OpenStudy (anonymous):

Simplify quantity x squared plus 8 times x all over x squared. x plus 8 all over x squared x plus 8 all over x 8x 8

OpenStudy (ranga):

\[\Large \frac{ x^2 + 8x }{ x^2 } = \frac{ x(x + 8) }{ x^2 } = \frac{ x + 8 }{ x }\]

OpenStudy (anonymous):

cool thanks can you help me with the next one

OpenStudy (ranga):

I can try.

OpenStudy (anonymous):

thanks i suck at math.. Simplify x squared minus 12 times x plus 20 all over x minus 10. x − 2 x minus 2 all over x minus 10 x minus 2 all over 10 x minus 10 all over x minus 2

OpenStudy (anonymous):

will it be C????????

OpenStudy (ranga):

No. You need to factor the numerator: x^2 - 12x + 20 Write it as: x^2 - 10x - 2x + 20 and try to factor it now.

OpenStudy (anonymous):

will be D then

OpenStudy (ranga):

See if I keep giving you answers then when you come across similar problems you won't be able to do it yourself. Try to factor the numerator. I gave you the first step: Write numerator as: x^2 - 10x - 2x + 20

OpenStudy (anonymous):



OpenStudy (anonymous):

is it is A

OpenStudy (ranga):

Yes! See you knew how to do it!